# and Reggio Emilia, Italy

### v2

@@ 4.1. Evaluation on Data Preparation Pipelines
4.2. Impact of Lazy Evaluation
{mdframed}
SummaryCuDF significantly outperforms other libraries when running the entire pipeline, while SparkSQL achieves remarkable results on all datasets. Lazy evaluation brings benefits mostly on large datasets, improving performance by 30% on average (up to almost 70% in some cases) compared to its eager counterpart.
Figure 5 reports for each library the runtime for the execution of the entire data preparation pipeline. For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL, and Polars), we also point out the difference compared to eager evaluation.
Considering the overall performance on the entire pipeline, CuDF consistently outperforms other libraries by effectively exploiting the power of GPU, with the only exception of the smallest dataset (Figure 5(a)). Pandas performs quite well on smaller datasets, but struggles on larger ones due to its eager approach (which requires computing intermediate results) causing higher memory consumption (Sinthong and Carey, 2021). Vaex and SparkSQL perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores for handling many string columns.
On one hand, it can be noticed how long logical plans can limit the optimizations introduced by lazy evaluation. Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 30% on Taxi and almost no changes in performance on Patrol. On the other hand, libraries like SparkPD leverage Pandas-specific optimizations (e.g., vectorized operations and memory efficiency) that translate into a 70% performance improvement. Further, Polars generally demonstrates good performance. In many cases, the eager API internally utilizes the lazy API before immediately collecting the result. Lazy execution in Polars leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 15%. However, it is important to notice that these optimizations are not always available, hence in some cases lazy operations may perform comparably to their eager counterpart (Figures 5(b) and 5(d)).
@@ 4.2. Impact of Lazy Evaluation
4.3. Scalability
@@
SummarySparkSQL is the best library for scaling to large datasets on a single machine. Indeed, for both Patrol and Taxi, it is the only one able to execute the entire pipelines with the laptop configuration. Polars follows closely, but requires a lot of resources in terms of RAM, exhibiting the worst scalability on machines with limited resources.
In this section, we assess how dataframe libraries scale by varying the size of the dataset and the configuration of the underlying machine. In particular, we take into account three different machine configurations, whose CPU and RAM are incrementally increased as described in Table 4 to reproduce typical configurations for laptops, workstations, and servers. Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi. Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi. Since CuDF scales according to GPU memory, it is not included in this experiment. If the dataset fits in the GPU, CuDF is able to efficiently handle it and complete the pipeline. However, this dependency does not make it a good candidate for dealing with very large datasets on machines with limited GPU resources.
Overall, SparkSQL clearly stands out as the best performer, being the only library capable of handling both datasets in their entirety on a laptop configuration. This capability is due to the use of disk spillover, which offloads data from RAM to disk storage, when the memory limit is reached during computations, ensuring continuous operation without exceeding memory capacity. On the other hand, SparkPD suffers instead from overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory errors on the laptop configuration for both Patrol and Taxi.
DataTable stands out on Patrol as one of the libraries that require the least resources, followed closely by ModinR and Pandas. Indeed, it is one of the libraries that consumes the least memory on Patrol; further, it does not require to split the dataframe into multiple pieces, which can lead to errors due to communication and information exchange.
ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart. The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to out-of-memory (OOM) issues more easily. In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach. This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards. Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized.
Despite its excellent performance in other experiments, scalability appears to be a significant weakness for Polars. This limitation arises from its in-memory execution model, which requires all data to be loaded into RAM. While this approach offers speed and efficiency for smaller datasets, it becomes a bottleneck when handling larger datasets, limiting its scalability.
Finally, Pandas confirms its well-known severe limitations about scalability, being the only library not able to complete the pipeline on Taxi even on a server configuration (Figure 6(c)).
@@
Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples. Laptop, Workstation, Server. See Table 4 for the detailed configurations.
@@
Pandas LP LP LP LP WS WS WS LP LP WS WS SV SV
SparkPD WS WS WS WS WS WS WS WS WS WS WS WS WS WS
SparkSQL LP LP LP LP LP LP LP LP LP LP LP LP LP LP
ModinD LP LP LP WS SV SV SV LP LP LP WS SV
ModinR LP LP LP LP WS WS WS LP LP LP LP WS WS WS
Polars LP LP LP WS WS WS WS LP LP WS WS WS SV SV
Vaex LP LP LP WS WS WS WS LP LP LP WS WS WS WS
Datatable LP LP LP LP LP WS WS LP LP LP LP WS WS WS
4.4. Performance on the TPC-H queries
{mdframed}
SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries.
TPC-H [^27] https://www.tpc.org/tpch/default5.asp is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000). It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data. Consistently with previous related work [^28] https://pola.rs/posts/benchmarks/ exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations [^29] https://github.com/Bodo-inc/Bodo-examples, replicating it for the other libraries, then use Bento for their seamless execution. Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark. We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100.
Overall, CuDF consistently achieves the best performance by leveraging GPU power. Among the CPU-only libraries, Polars is instead the clear winner. SparkSQL significantly outperforms Pandas on several queries, registering notable results especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%). Pandas is not the worst performer, but on most queries it does not achieve good results. Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query.
As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers. DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by their own benchmark [^30] https://h2oai.github.io/db-benchmark/) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09). Finally, Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
@@ 5. Related Work
To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks. In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources (Schmitt, 2020; Alexander, 2023; Pinner, 2023), which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations. The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations.
In literature, when a paper introduces a new dataframe library or dataframe-based tool, it usually includes a comparison against some dataframe libraries. It is the case for the AFrame data analysis package (Sinthong and Carey, 2019) (later extended to PolyFrame (Sinthong and Carey, 2021)), designed to efficiently manage large-scale dataframes using SQL++ queries. AFrame is evaluated against Pandas, Spark, and Pandas on Ray (i.e., ModinR) through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments. Similarly, Petersohn et al. (Petersohn et al., 2021) evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas. However, these comparisons share significant limitations, since data preparation is not their main focus and only a relatively small set of libraries and operations is considered.
Further, dataframe libraries are sometimes compared in data science benchmarks. For instance, as previously mentioned in Section 4.4, it is the case for TPC-H (Pöss and Floyd, 2000), which is used by Polars  [^31] https://pola.rs/posts/benchmarks/ to compare against Pandas, PySpark, Dask, Modin, and DuckDB (Raasveldt and Mühleisen, 2019). While the outcome of their analysis is consistent with our findings, their evaluation (still performed by selecting a scale factor of 10) only covers the first 7 queries out of the entire batch of 22 covered by our evaluation. Similarly, H2O (creator of DataTable) uses a set of queries to asses the scalability of popular database and dataframe-like systems [^32] https://h2oai.github.io/db-benchmark only limited to group by and join operations. Indeed, due to their different focus, such relational benchmarking systems notoriously offer very limited insights into real-world data preparation workflows executed on dataframes (Rehman and Elmore, 2022). Note that for producing a consistent a analysis we do not take into account tools that are not based on the dataframe data structure, as the aforementioned DuckDB or similar SQL-based solutions. For instance, Grizzly (Kläbe and Hagedorn, 2021) aims to address Pandas limitations in scalability by translating dataframes to SQL code, while PyFroid translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation.
@@
Sanzu (Watson et al., 2017) is designed to evaluate five popular data science frameworks and systems (R, Anaconda [^33] https://www.anaconda.com, Dask, PostgreSQL [^34] https://www.postgresql.org, and PySpark) on data processing and analysis tasks. Sanzu is composed of a micro-benchmark, to test basic operations in isolation on synthetic datasets of different sizes, and a macro-benchmark, to evaluate series of operations representing concrete data science scenarios on real-world datasets. FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations. Finally, Shanbhag et al. (Shanbhag and Chimalakonda, 2023) measure the energy consumption of 21 dataframe processing tasks on Pandas, Vaex, and Dask.
(a) Queries 1 to 11.
(b) Queries 12 to 22.
Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries.
@@ 6. Key Takeaways
Our evaluation confirmed the well-known limitations of Pandas. However, despite the abundance of alternative dataframe libraries designed to overtake it, there is no silver bullet to perform data preparation of tabular data with dataframes on a single machine. The right choice usually depends on many factors, such as the size and the features of the dataset at hand, the configuration of the underlying machine, and the kind of operations to carry out. We can help data practitioners to decide which library to use by answering the following questions.
@@
If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API. However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations [^35] https://pola-rs.github.io/polars/user-guide/misc/alternatives/.
@@
If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), DataTable stands out as the fastest library for I/O operations, while for other tasks it might not be that mature. Further, it shows the least compatibility with Pandas, requiring to implement many preparators manually. Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT and DC, also recording the best performance among all CPU-only libraries when running the entire pipeline.
@@
For small datasets (less than 500k rows), all of the libraries show similar average runtimes. Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, since their performance does not differ substantially—at least, not enough to learn a new API and replace entire parts of the pipeline (Pinner, 2023).
In the case of datasets of medium size (from 2 to 19 million rows) with dense numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed. On the other hand, for sparse datasets with many null values (e.g., Loan), Polars and SparkPD show good performance. Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.

### v3

@@ 4.3. Scalability
For this experiment, we selected the most computationally expensive pipeline among the three evaluated, which is the first one [^6] On average, this pipeline is approximately three times more computationally expensive than the others.. Since CuDF scales according to GPU memory, it is not included in this experiment. If the dataset fits in the GPU, CuDF is able to efficiently handle it and complete the pipeline. However, this dependency does not make it a good candidate for dealing with very large datasets on machines with limited GPU resources.
Figure 6 shows the performance of the libraries for running the entire pipeline on incremental samples of Taxi. SparkSQL clearly stands out as the best performer, being the only library capable of handling the full pipeline execution on laptop configuration (Figure 6(a)). This capability is attributed to its combination of lazy evaluation and disk spillover mechanisms (Apache Spark, 2024e)—optimizing execution plans and offloading data to disk when memory limits are reached during computations. SparkPD suffers instead from the overhead due to internal conversions between Pandas and Spark dataframes and increased JVM memory usage, causing out-of-memory (OOM) errors on the laptop configuration.
While Modin (both Dask and Ray) supports operations on datasets larger than available memory (Modin, 2024a), the efficacy of this feature is constrained by two primary factors. Firstly, it requires additional memory for managing internal operations, metadata, and intermediate results. Secondly, certain operations (e.g., apply lambda function), necessitate in-memory data access. DataTable shares this limitation as well, as a results these two libraries encounter OOM issues when processing datasets larger than 25% of the full dataset. Similarly, Vaex experiences OOM errors with datasets exceeding 15% of the full dataset size. This is due to the fact that in Vaex the output of some operations (e.g., groupby) is held entirely in memory, potentially exceeding available RAM (Vaex, 2024c).
Despite Polars excellent performance in other experiments, scalability appears its weakness—it reaches OOM with 4 million and 40 million rows in Figures 6(a) and  6(b), respectively. This limitation arises from its in-memory execution model, which requires all data to be loaded into RAM. While this approach offers speed and efficiency for smaller datasets, it becomes a bottleneck when handling larger datasets, limiting its scalability. Finally, Pandas confirms its well-known severe limitations about scalability, being the only library not able to complete the pipeline on Taxi even on a server configuration (Figure 6(c)).
To provide a comprehensive overview of how libraries scale across datasets with different characteristics, Table 5 outlines instead the minimum configuration required by each library to successfully run the full pipeline on progressively larger samples of Patrol and Taxi.
SparkSQL is the only library able to run the pipelines with the laptop configuration on the entire datasets—thanks to lazy evaluation and disk spillover. The second best in scalability is DataTable, which excels on Patrol requiring minimal resources through efficient memory mapping on disk.
@@
ModinR demonstrates better scalability across all machine configurations compared to its Dask counterpart. The Dask engine uses a centralized scheduler to distribute data across multiple cores, which results in higher memory consumption and leads to OOM issues more easily (Dask, 2024a). In contrast, Ray can scale far beyond Dask due to its distributed task scheduling scheme, specifically employing a distributed bottom-up scheduling approach (Ray, 2024a). This optimizes resource utilization and minimizes overhead by initiating tasks at the lowest level of computation and aggregating results upwards. Notably, Modin was originally built to work on top of Ray, making this integration more mature and optimized (Modin, 2024a).
(a) Queries 1 to 11.
(b) Queries 12 to 22.
Figure 7. Performance of the dataframe libraries on the TPC-H 10GB queries (lower is better).
4.4. Performance on the TPC-H Queries
@@
SummaryCuDF consistently achieves the best performance across all queries, while Polars significantly outperforms other CPU-only dataframe libraries. TPC-H (TPC-H, 2024) is a decision support benchmark consisting of a suite of business-oriented ad-hoc queries and concurrent data modifications (Pöss and Floyd, 2000). It is widely adopted to compare end-to-end database systems (Dreseler et al., 2020), allowing to comprehensively evaluate their performance and efficiency through the execution of complex queries on large volumes of data. Consistently with previous related work (Polars, 2024f) exploiting TPC-H queries to assess the performance of a subset of dataframe libraries (see Section 5 for more details), we also adopt this benchmark to provide further support to the outcomes of the comparative analysis presented above.
In particular, we take as a reference a publicly available translation of the TPC-H queries into Pandas API operations (Bodo, 2024), replicating it for the other libraries, then use Bento for their seamless execution. Figure 7 illustrates the performance of each library on the 22 queries of the TPC-H 10GB benchmark. We select a scale factor of 10, because it represents the largest dataset that can be processed by the 40 GB RAM of our GPU, which therefore would not be able to handle a scale factor of 100. Moreover, we included DuckDB (Raasveldt and Mühleisen, 2019), being a popular in-process SQL analytical database management system. DuckDB can execute parallel queries directly on Pandas DataFrames, Parquet/CSV files, or Arrow tables without a separate import step, and can write results back to these formats. This interoperability allows seamless integration with other data science libraries. We executed the TPC-H queries on DuckDB using the provided implementation in the Polars benchmark (Polars, 2024f). Notice that DuckDB does not provide Pandas API and only supports SQL. Thus, by expressing the pipelines as SQL queries, it can take advantage of all the well-known query optimization and execution strategies typical of relational database management systems. For this reason, we do not compare it with the other libraries throughout the paper, but only report its performance with TPC-H to provide a valuable reference point w.r.t. OLAP database management systems.
Overall, CuDF consistently achieves the best performance by leveraging GPU power. Among the CPU-only libraries, Polars is instead the clear winner. This confirms what is observed in the data preparation pipelines, for EDA and DT. SparkSQL  outperforms Pandas on several queries, registering notable results, especially on q06, where it is even faster than Polars (note that this query selects line items shipped within a year interval, reducing the dataset by 85%). By leveraging lazy evaluation, SparkSQL significantly reduces the amount of data processed early in the query execution—this is one of the key reasons why it improves performance in data preparation pipelines by approximately 60%. Pandas is never the worst performer, but on most queries, it is among the slowest. Modin (which registers the same results with both engines, Dask and Ray) shows diverse performance instead, strongly dependent on the query, confirming the inconsistent performance observed across the different data preparation stages.
As mentioned above, SparkPD suffers from the additional latency due to the translation of Pandas operations into Spark execution plans, making it one of the worst performers. DataTable is one of the worst performers too, since it is slow in grouping operations (as also testified by the execution of group preparator in DT stage and their own benchmark (H2O, 2021)) and its API only supports joins on columns with unique values, requiring therefore to switch to the default Pandas API (e.g., q09). Vaex is by far the worst performer overall, as it is significantly slow in grouping operations (as also pointed out in the previous sections) and lacks support for multi-column joins.
@@
Table 5. Minimum machine configuration for running the entire pipeline on incremental dataset samples.
@@
Pandas
SparkPD
SparkSQL
ModinD
ModinR
Polars
Vaex
DataTable
III Laptop   III Workstation   III Server   OOM
After CuDF, the second best performers are DuckDB and Polars. They excel in different aspects of the benchmark, with neither consistently outperforming the other across all queries—they adopt two different approaches for query optimizations.
@@ 5. Related Work
To the best of our knowledge, our paper presents the first rigorous and extensive experimental comparison of existing dataframe libraries on data preparation tasks. In fact, public wisdom about such libraries is mostly scattered across several not peer-reviewed sources (Schmitt, 2020; Alexander, 2023; Pinner, 2023), which generally compare the performance of few libraries (e.g., Pandas vs PySpark vs Polars) on a handful of operations. The major claims by such analysis are mostly confirmed by our evaluation; nevertheless, despite often providing some useful insights, they only offer a partial and very fragmented knowledge, far from a complete and detailed overview covering all libraries and operations. In the literature, only partial comparisons have been performed among libraries, which we list in the following.
AFrame (Sinthong and Carey, 2019, 2021) reports an evaluation against Pandas, Spark, and ModinR through a micro-benchmark using the synthetic Wisconsin benchmark dataset (DeWitt, 1993) to assess acceleration and scalability across distributed environments. Petersohn et al. (Petersohn et al., 2021) evaluate the benefits of Modin by comparing against Dask, Koalas, and Pandas. Grizzly (Kläbe and Hagedorn, 2021) evaluates against Modin and Pandas, while translating dataframe pipelines to SQL. Shanbhag et al. (Shanbhag and Chimalakonda, 2023) compare Pandas, Vaex and Dask on a set of single operations to gain insights into their energy consumption.
@@
Dataframe libraries have been compared also in data science benchmarks. For instance, Polars (Polars, 2024f) compares against Pandas, PySpark, Dask, Modin, and DuckDB on TPC-H (Pöss and Floyd, 2000). The outcome of their evaluation is consistent with our findings (we adopt the same scale factor for TPC-H), but their evaluation only covers the first 7 queries out of 22. Similarly, DataTable (H2O, 2021) uses a set of queries to asses the scalability of popular database and dataframe-like systems only limited to group by and join operations. Sanzu (Watson et al., 2017) proposes a benchmark designed to evaluate five popular data science frameworks and systems (R, Anaconda (Anaconda, 2024), Dask, PostgreSQL (PostgreSQL, 2024), and PySpark) on data processing and analysis tasks. It comprises both a micro-benchmark for testing basic operations in isolation and a macro-benchmark for evaluating series of operations representing concrete data science scenarios on real-world datasets. Similarly, FuzzyData (Rehman and Elmore, 2022) is a workflow generation system that allows to compare dataframe-based APIs on workflows composed of a small subset of operations.
Other works have explored the problem of optimizing/re-writing dataframe pipelines, reporting comparative evaluations of popular libraries. For instance, PyFroid (Emani et al., 2024) translates from the Pandas API to the DuckDB API to efficiently scale Pandas workloads on a commodity workstation, using top-voted Kaggle notebooks for the evaluation. Dataprep (Peng et al., 2021) is a task-centric EDA tool that can use different dataframe libraries as back-end engines, including Modin, PySpark, and Dask. Dias (Baziotis et al., 2024) is a system for dynamically rewriting Pandas code to optimize performance. All of these studies offer valuable insights into dataframe library performance across some real-world scenarios, but they do not provide a comprehensive evaluation on data preparation tasks with all dataframe libraries considered here.
@@ 6. Key Takeaways
Our evaluation confirmed the well-known limitations of Pandas. However, despite the abundance of dataframe libraries designed to overtake it, there is no silver bullet to perform the preparation of tabular data with dataframes on a single machine.
The relative performance of the libraries varies significantly among the variegate datasets and stages of the data preparation pipelines that we considered. None of the tools emerges as the clear winner for all considered scenarios. Nevertheless, certain characteristics of the dataset and the pipeline can help narrow the user’s choice of the tool. Below, we list three questions to assist in this selection.
@@
If a GPU with enough memory to fit the data is available, CuDF appears to consistently yield the best overall performance and has quasi-complete compatibility with the Pandas API [^7] This assumption is based on a single-machine scenario. It is important to note that while GPUs can provide substantial performance advantages, comparing GPUs and CPUs requires careful consideration of factors like costs (e.g., pay-per-use fees in a cloud environment) and the growing availability of high-core-count CPUs.. However, CuDF lacks of an optimizer; thus, libraries like Polars or SparkSQL might be faster when running the entire pipeline, avoiding unnecessary materializations (Polars, 2024b)
@@
If a user is only interested in an individual stage, i.e., input/output (I/O), explorative data analysis (EDA), data transformation (DT), or data cleaning (DC), CuDF and Polars stand out as the fastest libraries for reading and writing operation respectively. Polars emerges as the best performer for EDA, while SparkSQL represents a consistent solution for DT. Moreover, for DC, Vaex records the best performance among all CPU-only libraries when running the entire pipeline on larger datasets.
@@
For small datasets (less than 500k rows), all of the libraries show similar average runtimes. Even if Polars is often recommended for such a scenario, and indeed registers the best overall performance if memory resources are not limited, Pandas might still be the most reasonable choice, as their performance does not differ substantially—at least, not enough to replace entire parts of the pipeline (Pinner, 2023).
In the case of datasets of medium size (from 2 to 19 million rows) primarily composed of numeric columns (e.g., 5% Taxi sample), Vaex offers a robust solution if the pipeline involves many column-wise operations and Pandas compatibility is not needed. On the other hand, for datasets with a high proportion of missing values (e.g., Loan, where more than 30% of values are missing), ModinR and PySpark (both version) show good performance. Finally, for datasets with many string columns (e.g., 25% Patrol), SparkSQL appears to be the best choice.